# Unstable entropy for Anosov diffeomorphisms on the $3$-torus

### v1

@@ 1.1. Anosov diffeomorphisms with a partially hyperbolic splitting
Let $f\in\mathrm{Diffeo}^{\infty}(\mathbb{T}^{3})$ be a smooth diffeomorphism of $\mathbb{T}^{3}$. Suppose that $f$ is Anosov with splitting $T(\mathbb{T}^{3})=E_{s}\oplus E_{cu}$ where $\mathrm{dim}(E_{s})=1$ and $\mathrm{dim}(E_{cu})=2$. Suppose in addition that the unstable bundle $E_{cu}$ of $f$ splits as $E_{cu}=E_{c}\oplus E_{u}$ where $E_{c}$ is dominated by $E_{u}$, see §2 for the precise definitions. In other words, $f$ is a (strongly) partially hyperbolic diffeomorphism with splitting $T(\mathbb{T}^{3})=E_{u}\oplus E_{c}\oplus E_{s}$. We will call $E_{c}$ the center direction and $E_{u}$ the strong unstable direction of $f$. We further suppose that $f$ is orientation preserving [^1] This assumption is made to simplify the exposition. If $\mathrm{det}(f)<0$, one can consider $f\circ f$ instead. . We will denote by $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ the set of such diffeomorphisms.
@@ 1.2. Unstable entropy
For a partially hyperbolic diffeomorphism $f$ and an invariant probability measure $\mu$, Hu, Hua and Wu [24] defined the notion of metric entropy of $\mu$ with respect to the strong unstable foliation $E_{u}$, which we denote by $h^{u}_{\mu}(f)$. Another important invariant associated to $E_{u}$ is the unstable topological entropy $h^{u}_{\mathrm{top}}(f)$; an invariant introduced by Saghin and Xia [36] which captures the exponential growth rate of the volume of balls in the strong unstable manifold. The two notions of entropies are linked by a variational principle [24, Theorem D],
@@
where $\mathcal{P}_{f}(\mathbb{T}^{3})$ denotes the set of $f$-invariant probability measure on $\mathbb{T}^{3}$, see §2.2 for a quick review of these notions. A probability measure which attains the supremum is called a measure of maximal $u$-entropy or $u$-MME. Their set will be denoted by $\mathcal{M}^{u}(f)$. Hu,Wu and Zhu [25, Theorem B] showed that $\mathcal{M}^{u}(f)$ is always non-empty, convex, compact and that its extreme points are precisely the ergodic measures of maximal $u$-entropy. In this paper, we compute $h^{u}_{\mathrm{top}}(f)$ and study some properties of $\mathcal{M}^{u}(f)$. To state our results, let us define the center Jacobian,
@@
In particular, as an equilibrium state for an Anosov map, $\mu_{J^{c}_{f}}$ satisfies exponential mixing. Moreover, $\mu_{J^{c}_{f}}$ can be expressed via Bowen’s formula in terms of periodic points, see for instance [21, Theorem 20.3.7].
In a recent survey [38], Tahzibi asked whether for an Anosov diffeomorphism which is partially hyperbolic with expanding center, a measure of maximal $u$-entropy is necessarily a measure of maximal entropy. We will say that $E_{s}$ and $E_{u}$ are jointly integrable if there is a 2 dimensional invariant foliation tangent to $E_{s}\oplus E_{u}$. Using a criteria due to Gan and Shi [17], we show the following result.
@@
For $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, joint integrability of $E_{u}$ and $E_{s}$ is equivalent to $f$ being not accessible, see [42]. Accessibilty is $C^{r}$-dense ($r\geq 1$) [14, 7] and $C^{1}$-open [13] in $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. We thus answer negatively to [38, Question 2] for an open and dense set of diffeomorphisms $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$.
@@ 1.3. Margulis system of measures
Next, we study more precisely the structure of $\mathcal{M}^{u}(f)$. To this end, we will construct systems of measures $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ (resp. $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$) on the strong unstable (resp. center stable) manifolds. The study of measures of maximal entropy (and more generally equilibrium states) via system of leaf measures has a rich and long history, see the introductions of [12, 27] for an account of the existing literature.
@@
Theorem 2 should be compared to the construction of the measure of maximal entropy via systems of leaf measures. In the hyperbolic case, this is due to Sinai for maps [37] and to Margulis for flows [31]. Recently, different works have extended the construction of leaf measures to some partially hyperbolic settings, see §1.5 for a comparison with existing works. The Margulis system $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ can be used to characterize the measures of maximal unstable entropy.
@@
The author does not know if $\mu_{J^{c}_{f}}$ is the unique measure of maximal $u$-entropy or not. We note that because of the center direction being expanding, the usual Hopf argument does not seem to give the uniqueness as in [12, 9, 39].
@@ 1.4. Pollicott-Ruelle resonances
Here, $\mathrm{WF}(u)\subset T^{*}\mathbb{T}^{3}\setminus\{0\}$ denotes the wavefront set of $u$, see [23, Chapter VIII]. In this case, the current $u$ is called a resonant state associated to the resonance $\lambda$. We have a dual notion of co-resonant state:
@@
We say that $\lambda\in\mathbb{R}_{+}$ is the first resonance for the action on $k$-forms if $\lambda\in\mathrm{Res}_{k}(f)$ and $\mathrm{Res}_{k}(f)\subset\{z\in\mathbb{C},|z|\leq\lambda\}.$ For $k=0$ (resp. $k=1$), it is known that the first resonance is given by $1$ (resp. $e^{h_{\mathrm{top}}(f)}$) and that the associated (co)-resonant states can be used to reconstruct the SRB measure (resp. measure of maximal entropy), see [2, 18] for Anosov maps and [8, 27] for Anosov flows. For $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, it is therefore natural to ask what is the first resonance for the action on $2$-forms and whether the corresponding (co)-resonant states have dynamical meaning or not.
@@
where $g|_{\mathcal{N}}$ denotes the restriction of a current $g$ to a submanifold $\mathcal{N}.$222see for instance [23, Corollary 8.2.7]. Finally, the trace of the spectral projector at the resonance $\lambda=e^{h^{u}_{\mathrm{top}}(f)}$ is given by $\mu_{J^{c}_{f}}$.
@@
The restrictions in (1.8) are well defined by wavefront set conditions. In the case of the co-resonant state, we show in fact the stronger inclusion $\mathrm{WF}(\nu)\subset E_{cs}^{*}\subset E_{s}^{*}$, see (5.13), since the condition in (1.7) is a priori not enough to restrict a co-resonant state to an unstable manifold $\mathcal{W}^{u}(x)$.
@@ 1.5.1. Unstable entropy
The topological unstable entropy was shown to be upper semi-continuous by Yang [41] and continuous in our setting by Wu [40]. That $h^{u}_{\mathrm{top}}(f)$ is constant in a neighborhood of a linear automorphism $A$ follows from the work of Hua, Saghin and Xia [26]. Here, we extend this observation to all Anosov maps $f$ isotopic to $A$. The facts that the unstable topological entropy coincides with the pressure of $J^{c}_{f}$ and that $\mu_{J^{c}_{f}}\in\mathcal{M}^{u}(f)$ seem to be new, to the best of the author’s knowledge.
@@ 1.5.2. Systems of Margulis measures
Recently, different works have constructed systems of leaf measures for some classes of partially hyperbolic systems in order to study measures of maximal entropy, and more generally equilibrium states. Most relevant to our paper are the constructions of Climenhaga, Pesin and Zelerowicz [12], Parmenter and Pollicott [32], Carrasco and Rodriguez-Hertz [10], Buzzi, Fisher and Tazhibi [9] and Ures, Viana, Yang and Yang [39].
We emphasize that Theorem 2 does not follow from any of the previously cited works. Indeed, for $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, since the center direction $E_{c}$ is uniformly expanded, $f$ does not satisfy the Lyapunov stability condition of [12, 32], it is not a center isometry as required in [10], nor it is a flow type diffeomorphism as in [9]. Although any $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ is topologically conjugated to an Anosov automorphism $A$ of $\mathbb{T}^{3}$, it does not necessarily factor over Anosov as in [39]. Indeed, as shown in [42], the conjugacy maps the strong unstable foliation of $f$ to that of $A$ if and only if $E_{u}$ and $E_{s}$ are jointly integrable. As stated in the introduction, this condition fails for a dense and open subset of $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. For such $f$, the Margulis system $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ cannot be obtained by simply pulling back by the conjugacy the Lebesgue measure of $\mathcal{W}^{u}_{A}$, where $\mathcal{W}^{u}_{A}$ is the strong unstable manifold of $A$. When $f$ factor over Anosov, Corollary 1.2 does follow from [39, Theorem B] but since $f$ does not have c-mostly contracting center, [39, Theorem C] does not apply and the uniqueness of the $u$-MME is open even in this case. Actually, the author does not even know if a linear automorphism $A\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ has a unique $u$-MME or not.
To the best of the author’s knowledge, Theorem 2 is the first general construction of Margulis systems for a class of partially hyperbolic systems with uniformly expanding center. From a technical point of view, this makes showing the $\mathrm{Hol}^{cs}$-invariance (1.4) of $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ much harder due to the lack of regularity of center stable holonomies. To overcome this difficulty, we combine dynamical ideas inspired by [9, 10, 12] with functional techniques as in [27, 28], see the next subsection for an outline of the argument. We note that Alvarez, Leguil, Obata and Santiago studied $u$-Gibbs measures for $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ [^3] Their techniques actually allow them to study $C^{2}$ diffeomorphisms rather than smooth ones. in [1]. The lack of regularity of center stable holonomies is also aknowledged as a source of technical difficulty in their work, see [1, §2.2.3], although they use different techniques to overcome it.
@@ 1.5.3. Pollicott-Ruelle resonances
Theorem 3 seems to be the first result on the first resonance for Anosov maps for $k\neq 0,d_{s}$ where $d_{s}=\mathrm{dim}(E_{s})$. The close link between (co)-resonant states for $k=0,d_{s}$ and systems of leaf measures for Anosov diffeomorphism is well understood [18]. This was used by the author in [27] to reprove classical facts on the thermodynamical formalism for Anosov flows.
Recently, the author combined the functional approach for general Anosov actions obtained by Guedes Bonthonneau, Guillarmou, Hilgert, and Weich [19, 20] and the work of Carrasco and Rodriguez-Hertz [10] to study the measure of maximal entropy for minimal Anosov actions [28]. The present work can be seen as another instance of partially hyperbolic systems for which certain important dynamical invariant and measure can be reconstructed from the spectral approach.
We note that the functional approach was also used to construct physical and SRB measures for some partially hyperbolic systems, see [11] and the references therein.
@@ 1.6. Structure of the paper
In §3, we construct the unstable system of Margulis measures $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ using a compactness argument inspired by [9, 10]. The compactness used in the argument follows from a precise volume bound (3.4) which is a consequence of the work of Potrie [34]. We emphasize that since the complementary foliation $\mathcal{W}^{cs}$ is not contracting, this construction does not suffice to obtain the holonomy invariance (1.4) of $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ as in [9, 10]. Nevertheless, we show that $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ is measurable which allows us to deduce Corollary 1.2 using an argument of Tahzibi [38].
In §4, we recall some facts about Pollicott-Ruelle resonances and show Theorem 3 except for (1.8), using the functional approach. The argument is shorter than in [27, 28] since $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ is conjugated to a linear model, which we can use to simplify some proofs. We also deduce Theorem 1 and Corollary 1.1.
@@
  • we show using the functional approach that $\mu_{J^{c}_{f}}$ can be obtained as the limit of pullbacks of $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$, see Proposition 5.3. This should be compared with [12, Theorem 4.7 (1)] and [32, Theorem 1.1] although our techniques are quite different;
  • we deduce that $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$ is equal to the $cs$-conditionals of $\mu_{J^{c}_{f}}$ mimicking the dynamical argument of [12, Lemma 8.1];
  • using an adaptation of [12, Lemma 2.12], the local product structure of $\mu_{J^{c}_{f}}$ and Corollary 1.2, we deduce the holonomy invariance (1.4) of $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ from the holonomy invariance of $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$. This last argument seems to be the main novelty compared to previous works and might be of independent interest;
  • finally, we show (1.8) for $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ by adapting [27, Lemma 3.2].

### v3

@@ 1.1. Anosov diffeomorphisms with a partially hyperbolic splitting
Let $f\in\mathrm{Diffeo}^{\infty}(\mathbb{T}^{3})$ be a smooth diffeomorphism of $\mathbb{T}^{3}$. Suppose that $f$ is Anosov with splitting $T(\mathbb{T}^{3})=E_{s}\oplus E_{cu}$ where $\mathrm{dim}(E_{s})=1$ and $\mathrm{dim}(E_{cu})=2$. Suppose in addition that the unstable bundle $E_{cu}$ of $f$ splits as $E_{cu}=E_{c}\oplus E_{u}$ where $E_{c}$ is dominated by $E_{u}$, see §2 for the precise definitions. In other words, $f$ is a (strongly) partially hyperbolic diffeomorphism with splitting $T(\mathbb{T}^{3})=E_{u}\oplus E_{c}\oplus E_{s}$. We will call $E_{c}$ the center direction and $E_{u}$ the strong unstable direction of $f$. We further suppose that $f$ is transitive and orientation preserving [^1] This assumption is made to simplify the exposition. If $\mathrm{det}(f)<0$, one can consider $f\circ f$ instead. . Let $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ be the set of such diffeomorphisms.
@@ 1.2. Unstable entropy
For a partially hyperbolic diffeomorphism $f$ and an invariant probability measure $\mu$, Hu, Hua and Wu [undefw] defined the notion of metric entropy of $\mu$ with respect to the strong unstable foliation $E_{u}$, which we denote by $h^{u}_{\mu}(f)$. Another important invariant associated to $E_{u}$ is the unstable topological entropy $h^{u}_{\mathrm{top}}(f)$; an invariant introduced by Saghin and Xia [undefak] which captures the exponential growth rate of the volume of balls in the strong unstable manifold. The two notions of entropies are linked by a variational principle [undefw, Theorem D],
@@
where $\mathcal{P}_{f}(\mathbb{T}^{3})$ denotes the set of $f$-invariant probability measure on $\mathbb{T}^{3}$, see §2.2 for a quick review of these notions. A probability measure which attains the supremum is called a measure of maximal $u$-entropy or $u$-MME. Their set will be denoted by $\mathcal{M}^{u}(f)$. Hu,Wu and Zhu [undefx, Theorem B] showed that $\mathcal{M}^{u}(f)$ is always non-empty, convex, compact and that its extreme points are precisely the ergodic measures of maximal $u$-entropy. In this paper, we compute $h^{u}_{\mathrm{top}}(f)$ and study some properties of $\mathcal{M}^{u}(f)$. To state our results, let us define the center Jacobian,
@@
In particular, as an equilibrium state for an Anosov map, $\mu_{J^{c}_{f}}$ satisfies exponential mixing. Moreover, $\mu_{J^{c}_{f}}$ can be expressed via Bowen’s formula in terms of periodic points, see for instance [undeft, Theorem 20.3.7].
In a recent survey [undefam], Tahzibi asked whether for an Anosov diffeomorphism which is partially hyperbolic with expanding center, a measure of maximal $u$-entropy is necessarily a measure of maximal entropy. We will say that $E_{s}$ and $E_{u}$ are jointly integrable if there is a 2 dimensional invariant foliation tangent to $E_{s}\oplus E_{u}$. Using a criteria due to Gan and Shi [undefp], we show the following result.
@@
For $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, joint integrability of $E_{u}$ and $E_{s}$ is equivalent to $f$ being not accessible, see [undefaq]. Accessibilty is $C^{r}$-dense ($r\geq 1$) [undefm, undeff] and $C^{1}$-open [undefl] in $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. We thus answer negatively to [undefam, Question 2] for an open and dense set of diffeomorphisms $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$.
@@ 1.3. Margulis system of measures
Next, we study more precisely the structure of $\mathcal{M}^{u}(f)$. To this end, we will construct systems of measures $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ (resp. $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$) on the strong unstable (resp. center stable) manifolds. The study of measures of maximal entropy (and more generally equilibrium states) via system of leaf measures has a rich and long history, see the introductions of [undefk, undefz] for an account of the existing literature.
@@
Theorem 2 should be compared to the construction of the measure of maximal entropy via systems of leaf measures. In the hyperbolic case, this is due to Sinai for maps [undefal] and to Margulis for flows [undefad]. Recently, different works have extended the construction of leaf measures to some partially hyperbolic settings, see §1.5 for a comparison with existing works. The Margulis system $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ can be used to characterize the measures of maximal unstable entropy.
@@
The author does not know if $\mu_{J^{c}_{f}}$ is the unique measure of maximal $u$-entropy or not. We note that because of the center direction being expanding, the usual Hopf argument does not seem to give the uniqueness as in [undefk, undefh, undefan].
@@ 1.4. Pollicott-Ruelle resonances
Here, $\mathrm{WF}(u)\subset T^{*}\mathbb{T}^{3}\setminus\{0\}$ denotes the wavefront set of $u$, see [undefv, Chapter VIII]. In this case, the current $u$ is called a resonant state associated to the resonance $\lambda$. We have a dual notion of co-resonant state:
@@
We say that $\lambda\in\mathbb{R}_{+}$ is the first resonance for the action on $k$-forms if $\lambda\in\mathrm{Res}_{k}(f)$ and $\mathrm{Res}_{k}(f)\subset\{z\in\mathbb{C},|z|\leq\lambda\}.$ For $k=0$ (resp. $k=1$), it is known that the first resonance is given by $1$ (resp. $e^{h_{\mathrm{top}}(f)}$) and that the associated (co)-resonant states can be used to reconstruct the SRB measure (resp. measure of maximal entropy), see [undefa, undefq] for Anosov maps and [undefg, undefz] for Anosov flows. For $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, it is therefore natural to ask what is the first resonance for the action on $2$-forms and whether the corresponding (co)-resonant states have dynamical meaning or not.
@@
where $g|_{\mathcal{N}}$ denotes the restriction of a current $g$ to a submanifold $\mathcal{N}.$222see for instance [undefv, Corollary 8.2.7]. Finally, the trace of the spectral projector at the resonance $\lambda=e^{h^{u}_{\mathrm{top}}(f)}$ is given by $\mu_{J^{c}_{f}}$.
@@
The restrictions in (1.8) are well defined by wavefront set conditions. In the case of the co-resonant state, we show in fact the stronger inclusion $\mathrm{WF}(\nu)\subset E_{cs}^{*}\subset E_{s}^{*}$, see (5.13), since the condition in (1.7) is a priori not enough to restrict a co-resonant state to a strong unstable manifold $\mathcal{W}^{u}(x)$.
@@ 1.5.1. Unstable entropy
The topological unstable entropy was shown to be upper semi-continuous by Yang [undefap] and continuous in our setting by Wu [undefao]. That $h^{u}_{\mathrm{top}}(f)$ is constant in a neighborhood of a linear automorphism $A$ follows from the work of Hua, Saghin and Xia [undefy]. Here, we extend this observation to all Anosov maps $f$ isotopic to $A$. The facts that the unstable topological entropy coincides with the pressure of $J^{c}_{f}$ and that $\mu_{J^{c}_{f}}\in\mathcal{M}^{u}(f)$ seem to be new, to the best of the author’s knowledge.
@@ 1.5.2. Systems of Margulis measures
Recently, different works have constructed systems of leaf measures for some classes of partially hyperbolic systems in order to study measures of maximal entropy, and more generally equilibrium states. Most relevant to our paper are the constructions of Climenhaga, Pesin and Zelerowicz [undefk], Parmenter and Pollicott [undefae], Carrasco and Rodriguez-Hertz [undefi], Buzzi, Fisher and Tazhibi [undefh] and Ures, Viana, Yang and Yang [undefan].
We emphasize that Theorem 2 does not follow from any of the previously cited works. Indeed, for $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, since the center direction $E_{c}$ is uniformly expanded, $f$ does not satisfy the Lyapunov stability condition of [undefk, undefae], it is not a center isometry as required in [undefi], nor it is a flow type diffeomorphism as in [undefh]. Although any $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ is topologically conjugated to an Anosov automorphism $A$ of $\mathbb{T}^{3}$, it does not necessarily factor over Anosov as in [undefan]. Indeed, as shown in [undefaq], the conjugacy maps the strong unstable foliation of $f$ to that of $A$ if and only if $E_{u}$ and $E_{s}$ are jointly integrable. As stated in the introduction, this condition fails for a dense and open subset of $\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$. For such $f$, the Margulis system $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ cannot be obtained by simply pulling back by the conjugacy the Lebesgue measure of $\mathcal{W}^{u}_{A}$, where $\mathcal{W}^{u}_{A}$ is the strong unstable manifold of $A$. When $f$ factor over Anosov, Corollary 1.2 does follow from [undefan, Theorem B] but since $f$ does not have c-mostly contracting center, [undefan, Theorem C] does not apply and the uniqueness of the $u$-MME is open even in this case. For a linear automorphism $A\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$, the foliations are linear and are in particular absolutely continuous. This means that using Corollary 1.2 and the standard Hopf argument, one deduces that $A$ has a unique $u$-MME, see for instance [undefh, Proposition 3.8].
To the best of the author’s knowledge, Theorem 2 is the first general construction of Margulis systems for a class of partially hyperbolic systems with uniformly expanding center. From a technical point of view, this makes showing the $\mathrm{Hol}^{cs}$-invariance (1.4) of $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ much harder due to the lack of regularity of center stable holonomies. To overcome this difficulty, we combine dynamical ideas inspired by [undefh, undefi, undefk] with functional techniques as in [undefz, undefaa], see the next subsection for an outline of the argument. We note that Alvarez, Leguil, Obata and Santiago studied $u$-Gibbs measures for $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ [^3] Their techniques actually allow them to study $C^{2}$ diffeomorphisms rather than smooth ones. in [undef]. The lack of regularity of center stable holonomies is also aknowledged as a source of technical difficulty in their work, see [undef, §2.2.3], although they use different techniques to overcome it.
@@ 1.5.3. Pollicott-Ruelle resonances
Theorem 3 seems to be the first result on the first resonance for Anosov maps for $k\neq 0,d_{s}$ where $d_{s}=\mathrm{dim}(E_{s})$. The close link between (co)-resonant states for $k=0,d_{s}$ and systems of leaf measures for Anosov diffeomorphism is well understood [undefq]. This was used by the author in [undefz] to reprove classical facts on the thermodynamical formalism for Anosov flows.
Recently, the author combined the functional approach for general Anosov actions obtained by Guedes Bonthonneau, Guillarmou, Hilgert, and Weich [undefr, undefs] and the work of Carrasco and Rodriguez-Hertz [undefi] to study the measure of maximal entropy for minimal Anosov actions [undefaa]. The present work can be seen as another instance of partially hyperbolic systems for which certain important dynamical invariant and measure can be reconstructed from the spectral approach.
We note that the functional approach was also used to construct physical and SRB measures for some partially hyperbolic systems, see [undefj] and the references therein.
@@ 1.6. Structure of the paper
In §3, we construct the unstable system of Margulis measures $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ using a compactness argument inspired by [undefh, undefi]. The compactness used in the argument follows from a precise volume bound (3.4) which is a consequence of the work of Potrie [undefag]. We emphasize that since the complementary foliation $\mathcal{W}^{cs}$ is not contracting, this construction does not suffice to obtain the holonomy invariance (1.4) of $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ as in [undefh, undefi]. Nevertheless, using a Carathéodory construction inspired by Climenhaga, Pesin and Zelerowicz [undefk], we show that $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ is measurable using which allows us to deduce Corollary 1.2 using an argument of Tahzibi [undefam].
In §4, we recall some facts about Pollicott-Ruelle resonances and show Theorem 3, except for (1.8), using the functional approach. The argument is shorter than in [undefz, undefaa] since $f\in\mathcal{A}_{+}^{\infty}(\mathbb{T}^{3})$ is conjugated to a linear model, which we can use to simplify some proofs. We also deduce Theorem 1 and Corollary 1.1.
@@
  • we show using the functional approach that $\mu_{J^{c}_{f}}$ can be obtained as the limit of pullbacks of $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$, see Proposition 5.3. This should be compared with [undefk, Theorem 4.7 (1)] and [undefae, Theorem 1.1] although our techniques are quite different;
  • we deduce that $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$ is equal to the $cs$-conditionals of $\mu_{J^{c}_{f}}$ mimicking the dynamical argument of [undefk, Lemma 8.1];
  • using an adaptation of [undefk, Lemma 2.12], the local product structure of $\mu_{J^{c}_{f}}$ and Corollary 1.2, we deduce the holonomy invariance (1.4) of $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ from the holonomy invariance of $(\mu^{cs}_{x})_{x\in\mathbb{T}^{3}}$. This last argument seems to be the main novelty compared to previous works and might be of independent interest;
  • finally, we show (1.8) for $(\mu^{u}_{x})_{x\in\mathbb{T}^{3}}$ by adapting [undefz, Lemma 3.2].
@@
The author would like to further thank Ali Tahzibi for noticing that the previous argument given in Remark 3 was incomplete and Jérôme Buzzi for making him aware of a gap in the previous proof of the measurability of the Margulis system in Proposition 3.4. These issues have now been resolved in the current version of the paper. Finally, he would like to thank Ali Tahzibi for pointing out to him that the $u$-MME is unique for linear automorphisms and Gabriel Rivière for pointing out the work [undefaj], which lead to Remark 2.